Suppose we have the following values for variables X (independent) and Y (dependent) _ _ Σ(X – X)(Y – Y) = -630 _ Σ(X – X)2 = 168 SSE = 1225 SSR = 2362.5 If n = 8, what is the value of the standard error for the simple regression model (round to three decimals)?
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

Standard error , s = sqrt(MSE)

But MSE = SSE/(n-m), where m = 1 ( number of predictors), and n = sample size = 8

MSE= 1225/7 = 175

So, s = sqrt(175) = 13.229

Hence, answer is : Standard error = 13.229